# EmotiCrafter: Text-to-Emotional-Image Generation based on Valence-Arousal Model

### v1

@@ 4.1 Comparisons
Baselines. We built baselines for comparison via two strategies: (1) inject the emotion features directly into the image generation modules in the SDXL such as UNet; (2) use the emotion features to change the input text prompt that controls the content of the generated image (similar to the proposed method).
@@
Qualitative Comparison. We evaluate the generated images based on three criteria: (1) the effectiveness of emotion embedding, (2) image-prompt similarity, and (3) the continuity of the image varies as V/A values change. Figure 4 illustrates that the baseline methods, Cross Attention, Time Embedding, and Textual Inversion, fail to reveal the continuous change of emotions. Notably, Textual Inversion tends to generate images with a persistent purple tint, which significantly limits its emotional expressiveness. Only our method and the GPT-4+SDXL successfully generate images that clearly reflect emotional variations. It also demonstrates that all methods generate content faithfully to the given prompts. For image continuity, we focus on comparing our method and GPT-4+SDXL, as they are the only two that successfully generate emotional variations. Our method shows enhanced continuity, indicating superior controllability over continuous V-A values. Figure 5 offers a more comprehensive comparison, revealing that our approach maintains continuity even under extreme V/A conditions. Conversely, GPT-4+SDXL displays noticeable discontinuities (e.g., in the V=3 column).
Quantitative Comparison. We compare our method against several baselines using the following metrics: (1) V/A-Error evaluates the absolute error between the predicted V/A values of the generated images and the input V/A values. (2) CLIPScore [11] assesses the similarity between the input text and the generated images. (3) CLIP-IQA [43] leverages a pre-trained CLIP model to evaluate image quality without requiring reference images. (4) LPIPS-Continuous utilizes the Learned Perceptual Image Patch Similarity [59] to measure the continuity of the change in image as V / A changes.
@@
Table 1: Quantitative results comparing emotion accuracy, prompt fidelity, and image quality across different baselines (132 prompts × 5 V/A values × 25 images = 3,300 images per method). Our method achieves the highest performance in emotion accuracy, while maintaining comparable results in prompt fidelity and image quality, with a minor decrease due to the fact that altering emotional content will inherently affect semantic alignment.
As shown in Table 1, our method achieves the best (lowest) V/A-Error on average. Cross Attention and Time Embedding achieve the best (highest) CLIP-IQA and CLIPScore respectively. However, these methods fail to generate emotional images, as shown in Figure 4. Our method exhibits a slight decrease in CLIPScore relative to the baselines. We believe this is due to the fact that altering emotional content will inherently affect semantic alignment. At the same time, CLIP-IQA value shows that our method could generate high-quality images. Our method demonstrates superior continuity when compared to that of GPT-4 + SDXL (Table 2).
@@
Table 2: The continuity comparison between our method and GPT-4+SDXL (baseline).
@@ 4.2 User Study
A-Ranking Consistency $\uparrow$ 0.759±0.273 0.165±0.379
@@
Emotion Consistency $\uparrow$ 4.215±0.715 3.525±1.065
@@
Table 3: User study. Our method outperformed the baseline in all aspects.
Experiment I. The first experiment tested whether the emotion represented by the generated images aligned with a human’s perception. To this end, two collections of images were prepared respectively for testing the perception of A and V values. Each collection consists of 20 sets of images with 10 sets generated by our method and the other 10 sets generated by the baseline. Each set consisted of 5 images with different A or V values whose orders were randomized. The participants were asked to reorder these 5 images respectively based on the perceived intensity of V or A. Kendall’s $\tau_{b}\in[-1,+1]$ was computed to assess the alignment between the order given by participants and the ground truth order, where $\tau_{b}=1$ indicated a full match of the orders. During the experiment, We also asked participants to use their perceived emotions to estimate and label the V and A values of each image. The absolute error between the results and the ground truth was calculated.
Experiment II. The second experiment was designed to test whether the generated images were able to reveal the contentious change of emotions (i.e., V-A values). To this end, we generated 20 sets of images (10 based on our method and the other 10 based on baseline). As shown in Figure 5, each set consists of 25 images generated based on V-A values gradually changed from -3 to +3. Participants were asked to rate each image set (using a 5-point Likert scale) regarding two aspects: (1) the consistency between the change of V-A values and the change of image content; (2) the smoothness of the change of the image content.
Analysis & Results. We conducted a Shapiro-Wilk test [38] to assess normality and applied the Wilcoxon Signed Rank test [46] to evaluate the significance of all the results. The results showed that our method outperformed the baseline in all aspects (Table 3) and showed a statistical significance regarding V/A Ranking, V/A Error, Consistency, and Smoothness (Figure 6).
Figure 6: Significance test. (a) V/A Ranking: Kendall’s $\tau$b, (b) V/A Error: absolute error, (c) Emotion consistency, and Emotion smoothness. Significance levels: *p<0.05, **p<0.01, ***p<0.001
Figure 7: Abalation Study. Images are generated from the prompt “An oil painting shows an astronaut.” As $\alpha$ increases, image-prompt similarity decreases, while emotional variations increase. The usage of $d(a,v)$ enhances the accuracy of emotional changes in the generated images.
@@ 4.3 Ablation Study
We conducted ablation studies to estimate the effectiveness of the proposed loss function.
Effectiveness of the Scaling Factor $\alpha$. We tested the changes of CLIPScore and V/A-Error while the scaling factor $\alpha$ varies. The results (Figure 8) showed that when $\alpha$ increased, both the CLIPScore and V/A-Error decreased (CLIPScore getting worse, V/A-Error getting better). This finding can also be observed from the examples shown in Figure 7. We set $\alpha=1.5$ as a trade-offer regarding the experiment results. However, users still can choose different $\alpha$ values to customize their output.
@@
Effectiveness of the Density Weighting $d(v,a)$. We compare our method with a variant that removed $d(v,a)$ from the training process. As shown in Table 4, incorporating $d(v,a)$ into the loss function positively affects both the CLIPScore and V/A-Error, which can also be easily found in Figure 7.
4.4 Discussion
Although our method achieved promising results, it still has some limitations that are worth mentioning: First, we observe that controlling image generation based on arousal is more challenging compared to valence. This aligns with prior research in visual emotion analysis, which has similarly reported that arousal is harder to predict, largely due to lower agreement among annotators when assessing arousal levels [26]; Second, our approach often generates images featuring human activities, even when such activities are not specified in the prompts. This tendency is likely attributable to the limited representation of non-human scenes in the training data and could be addressed by incorporating a greater variety of non-human scenarios into the dataset; Third, our method occasionally adjusts users’ input prompts to better align with the specified emotional prompts, resulting a slight shift of the prompt meaning, thus effects the image generation. We believe this issue could be addressed by adding a term for semantic preserving in the loss function.
5 Conclusion
In this paper, we introduce a new task, i.e., continuous emotional image content generation (C-EICG), and present EmotiCrafter, a novel method for generating emotionally expressive images based on continuous emotional values (Valence-Arousal). Specifically, we introduce an emotion-embedding network that integrates V-A values into textual features. Extensive experiments demonstrate the effectiveness of our approach, showing that it can reliably generate images that align with both user-provided prompts and specified emotions described by V-A values. We believe this work will catalyze further advancements in affective computing and image generation. To encourage future research, we will release our code and data, hoping it will inspire new developments in the field.

### v2

@@ 4 Evaluation
4.1 Generation Results
Figure 1 shows the proposed technique’s ability to achieve continuous and effective control over both emotion and content during image generation. Meanwhile, Figure 5 highlights four key capabilities: (a) emotion-content decoupling, where V-A values override emotional cues in the prompt, allowing typically positive concepts to be rendered with negative emotions; (b) compatibility with discrete emotion categories; (c) content-independent generation, where images generated from empty prompts and specified V-A values maintain emotional consistency without semantic constraints; and (d) fine-grained emotional control, demonstrated through V-A increments of 0.2, showcasing the model’s sensitivity to subtle emotional variations.
@@ 4.2 Comparisons
Baselines. We established baselines for comparison using two strategies: (1) directly injecting emotion features into the image generation modules of SDXL, such as UNet, and (2) modifying the input text prompt using emotion features to influence the generated image’s content, similar to the proposed method.
@@
Qualitative Comparison. We evaluate the generated images based on three criteria: (1) the effectiveness of emotion embedding, (2) image-prompt similarity, and (3) the continuity of emotional variations as V-A values change. Figure 6 demonstrates that the baseline methods—Cross Attention, Time Embedding, and Textual Inversion—tend to produce nearly identical outputs regardless of emotional variation. This is because the loss terms of these baselines primarily align low-level image features (e.g., SDXL’s latent space), which are highly correlated with prompt content but struggle to capture subtle emotional cues when both prompts and V-A values are provided. Notably, Textual Inversion often produces images with a persistent purple tint. Furthermore, all methods faithfully generate content aligned with the given prompts. However, for image continuity, we focus on comparing our method with GPT-4+SDXL, as they are the only two capable of generating distinct emotional variations. Figure 7 provides a more detailed comparison, showing that our method maintains smooth emotional transitions even under extreme V-A conditions, whereas GPT-4+SDXL introduces noticeable discontinuities (e.g., in the V=3 column).
Quantitative Comparison. We compare our method against several baselines using the following metrics: (1) V/A-Error evaluates the absolute error between the predicted V/A of the generated images and the input V/A. (2) CLIPScore [8] assesses the similarity between the input text and the generated images. (3) CLIP-IQA [41] leverages a pre-trained CLIP model to evaluate image quality without requiring reference images. (4) LPIPS-Continuous utilizes the Learned Perceptual Image Patch Similarity [56] to measure the continuity of the change in image as V/A changes.
@@
Table 1: Comparison on emotion accuracy, prompt fidelity, and image quality across different baselines, evaluated on 3,300 images per method (132 prompts × 5 V values × 5 A values). Our method achieves the highest performance in emotion accuracy while maintaining comparable results in prompt fidelity and image quality. The slight decrease in prompt fidelity is expected, as modifying emotional content affects semantic alignment.
As shown in Table 1, our method achieves the lowest (best) V/A-Error on average. While Cross Attention and Time Embedding achieve the highest CLIP-IQA and CLIPScore, respectively, these methods fail to generate emotionally expressive images (Figure 6).
Our method exhibits a slight decrease in CLIPScore compared to the baselines, which we attribute to the inherent trade-off between emotional modulation and strict semantic alignment. However, the high CLIP-IQA score indicates that our method produces high-quality images. Additionally, as shown in Table 2, our approach demonstrates superior continuity compared to GPT-4 + SDXL.
@@
Table 2: Continuity comparison between our method and baseline.
@@ 4.3 User Study
Study I A-Ranking Consistency $\uparrow$ 0.759±0.273 0.165±0.379
@@
Study II Emotion Consistency $\uparrow$ 4.215±0.715 3.525±1.065
@@
Table 3: Our method outperformed the baseline.
Study I. In the first experiment, we evaluated whether the generated images’ emotions align with human perception. Two image collections were prepared—one for Arousal (A) and one for Valence (V)—each comprising 20 sets (10 from our method and 10 from the baseline) of 5 randomized images with varying A or V values. Participants reordered the images by perceived intensity and estimated each image’s V and A values. We then computed Kendall’s $\tau_{b}$ (with $\tau_{b}=1$ indicating perfect alignment) to assess ordering accuracy and calculated the absolute error between the estimated and ground truth values.
Figure 8: User Study Results ( *p<0.05, **p<0.01, ***p<0.001)
Figure 9: Abalation Study. Images are generated from the prompt “An oil painting shows an astronaut.” As $\alpha$ increases, image-prompt similarity decreases, while emotional variations increase. The usage of $d(a,v)$ enhances the accuracy of emotional changes.
Study II. In the second experiment, we assessed whether the generated images could effectively reflect continuous emotional changes (i.e., V-A values). We generated 20 image sets (10 per method), each containing 25 images with V-A values varying gradually from -3 to +3 ( Figure 7). Participants rated each set on a 5-point Likert scale regarding (1) the alignment between V-A changes and image content, and (2) the smoothness of the content transition.
Analysis & Results. We conducted a Shapiro-Wilk test [35] to assess normality and applied the Wilcoxon Signed Rank test [44] to evaluate statistical significance. Our results indicate that our method outperformed the baseline across all metrics (Table 3), with statistically significant improvements in V/A Ranking, V/A Error, Consistency, and Smoothness (Figure 8).
@@ 4.4 Ablation Study
We performed ablation experiments to assess the contribution and effectiveness of the proposed loss function.
Effectiveness of the Scaling Factor $\alpha$. We evaluated how varying $\alpha$ affects CLIPScore and V/A-Error (Figure 10). As $\alpha$ increases, CLIPScore decreases (indicating reduced semantic alignment), while V/A-Error also decreases (indicating improved emotional accuracy). This trend is further illustrated by the examples in Figure 9. Based on these findings, we set $\alpha=1.5$ as an optimal trade-off, though users can adjust $\alpha$ to suit their specific needs.
@@
Effectiveness of the Density Weighting $d(v,a)$. We compare our full method with a variant that omits d(v,a) from the loss function. As shown in Table 4, including d(v,a) leads to improvements in both CLIPScore and V/A-Error. This positive effect is further illustrated in Figure 9.
5 Conclusion and Limitations
In this paper, we introduce continuous emotional image content generation (C-EICG) and present EmotiCrafter, a novel method that generates emotionally expressive images using continuous Valence-Arousal (V-A) values. Our emotion-embedding network integrates V-A values into textual features, and extensive experiments show that our approach reliably aligns images with both user prompts and specified emotions. We believe this work will advance affective computing and image generation, and we will release our code and data to foster further research.
Although our method achieved promising results, it still has some limitations need to be addressed in the future. First, controlling image generation based on arousal remains more challenging than controlling based on valence. This is consistent with prior research in visual emotion analysis, which has found that arousal is harder to predict due to lower inter-annotator agreement [21]. Second, our approach frequently generates images featuring human activities even when such activities are not mentioned in the prompts. This likely stems from the limited representation of non-human scenes in our training data and could be mitigated by incorporating a more diverse range of non-human scenarios. Third, our method occasionally modifies users’ input prompts to better align with the specified emotional prompts, resulting in a slight semantic shift that affects the generated images. We believe this issue could be addressed by adding a semantic preservation term to the loss function.
6 Acknowledgments
Nan Cao is the corresponding author. This work was supported by the National Key Research and Development Program of China (2023YFB3107100).